# Proxy-Tuning: Tailoring Multimodal Autoregressive Models for Subject-Driven Image Generation

### v1

@@ 1 Introduction
Recently, multimodal autoregressive (AR) models [5, 25, 15, 9, 29], based on next-token prediction and transformer architecture [27], have demonstrated remarkable capabilities across various multimodal tasks, including text-to-image (T2I) generation. These models have shown comparable performance to large diffusion models [21, 19, 7] in general image generation tasks. However, the application of AR models to subject-driven image generation, a task of significant user interest [22], remains relatively unexplored compared to the extensive research on diffusion models in this domain [22, 4, 1, 8, 34, 14, 32, 13, 31].
Subject-driven image generation [22, 8, 12] requires a model to learn the visual appearance of a specific subject from a small set of reference images and then generate novel images of that subject based on user prompts. While diffusion models [10, 21] have been successfully adapted for this task using methods like DreamBooth [22], our initial attempts to tailor AR models for subject-driven generation revealed significant challenges. We explore two primary approaches to adapt AR models for this task: parameter-efficient fine-tuning using LoRA [11] and end-to-end fine-tuning. In both cases, we bound the given subject to a pre-defined token (e.g., S*) for use during inference. However, our experiments reveal limitations in both approaches. LoRA-based fine-tuning struggles to capture the specific subject’s appearance accurately, while end-to-end fine-tuning leads to diminished text fidelity in the generated images and compromises the model’s original capabilities.
To address these limitations, we introduce Proxy-Tuning, a novel method that leverages diffusion models to enhance AR models’ capabilities in subject-specific image generation. Our approach involves first fine-tuning a relatively weaker diffusion model on the reference images of the specific subject, and then uses the text-image pairs generated by this fine-tuned diffusion model to supervise the fine-tuning of more powerful AR models. Surprisingly, we discover a striking weak-to-strong generalization [2] phenomenon: fine-tuned AR models consistently outperform their diffusion model supervisors in both subject fidelity and prompt adherence. This trend holds true across different combinations and architectures of teacher (diffusion) and student (AR) models. Further analysis reveals scenarios where AR models particularly excel, e.g., multi-subject compositions and contextual understanding.
@@ 2.1 Subject-Driven Image Generation
Subject-driven image generation is a task that focuses on generating images related to a specific subject, where the model learns from a few reference images of the subject and combines with given prompts to create novel images of that subject. Currently, the existing methods regarding subject-driven image generation are mainly divided into two categories. The first category is tuning-based methods [22, 4, 1, 8, 12], which require fine-tuning the model for each specific target. Specifically, DreamBooth [22] binds the given subject to a specially pre-defined token (e.g., S*) and introduces a class-specific prior preservation loss to preserve the original knowledge of the pretrained diffusion model. Disenbooth [4] proposes to enhance the disentanglement of the specific subject and background features during the generation process, aiming to improve the quality and flexibility of the generated images. NeTI [1] focuses on improving the neural network architectures used in subject-driven image generation to achieve better visual fidelity and semantic consistency. The another category is tuning-free methods [34, 14, 32, 33, 28, 16, 23, 30]. Among these methods, IP-Adapter [34] introduces a decoupled cross-attention mechanism to capture and fuse the image embedding input and text embedding input to guide the process of image generation. PhotoMaker [14] extracts image features by using a pretrained CLIP image encoder and projects the image features to the text latent space through a lightweight mapper. Eventually, these image features are concatenated with the text features and then input into the model. Similarly, ELITE [30] also extracts image features by using a pretrained encoder, and incorporates the image features into the text latent space and the cross-attention module respectively by means of global mapping and local mapping.
@@ 2.2 Multimodal Autoregressive Models
Multimodal autoregressive (AR) models [24, 25, 15, 29, 9, 5], based on next-token prediction and transformer architecture, have shown impressive abilities in various multimodal tasks, especially in the realm of text-to-image (T2I) generation. LlamaGen [24] explores the application of the ”next-token prediction” paradigm from large language models to visual generation. By scaling the widely used autoregressive Llama backbones, it attains excellent image generation performance. Similarly, Chameleon [25] also adopts the autoregressive generation approach, capable of transforming images and text into discrete tokens for processing, showing excellent performances in various tasks such as visual question answering and image generation. Based on Chameleon, Lumina-mGPT [15] enhances the training datasets and achieves flexible text-to-image generation and controllable generation, also capable of handling vision recognition tasks. Moreover, Emu3 [29] further includes the video datasets in the training datasets and trains the transformer model only with next-token prediction by encoding images, texts and videos into a discrete space and jointly training on multimodal mixed sequences.
@@
LoRA SDXL [19] 0.8002 0.3225 0.7272
End-to-End Lumina-mGPT [15] 0.6974 0.2652 0.5338
LoRA Lumina-mGPT [15] 0.6752 0.2956 0.5088
Table 1: Quantitative evaluation of DreamBooth [22] subject-driven fine-tuning on AR model. Details about the metrics are provided in Section 5.1.
Figure 2: Framework of our Proxy-Tuning. Provided with a limited number of images of a particular target, we initially conduct subject-tuning (i.e., DreamBooth) on the diffusion model to acquire the specific target’s features. Subsequently, the diffusion model is employed as a supervisor to supervise the fine-tuning procedure of the AR model. Direct subject-tuning of the AR model results in an unsatisfactory target appearance, and the use of data augmentations in subject-tuning manifests a deficiency in semantic editability (as discussed in Section 5.6). Conversely, Proxy-Tuning effectively captures the target appearance and simultaneously showcases excellent semantic editability.
3 Analysis: Subject-Driven AR Tuning
To explore subject-driven image generation capabilities in AR models, we initially explored the settings in DreamBooth [22] of fine-tuning models with few subject images. Our investigation focused on two primary scenarios: end-to-end fine-tuning of the entire model and parameter-efficient tuning using LoRA [11]. Following the DreamBooth paradigm, we associate the target subject with a designated token (e.g., ”S*”) for inference purposes.
@@ 3 Analysis: Subject-Driven AR Tuning
Notably, these challenges are less prominent in diffusion models under similar few-shot subject-driven settings [22]. This disparity can be attributed to the inherent characteristics of AR models: they generate images through sequential token predictions where each token heavily depends on previous ones. This autoregressive nature makes them particularly sensitive to parameter updates when trained on very few subject images, as even slight perturbation in token prediction can propagate and amplify throughout the generation sequence. In contrast, diffusion models’ parallel denoising process demonstrates greater robustness to few-shot fine-tuning, better preserving both subject fidelity and semantic consistency.
@@ 4 Method: Proxy-Tuning
Notably, we discover a compelling weak-to-strong phenomenon in image generation, similar to observations in natural language processing [2]. Our findings show that AR models, when trained under the supervision of a weak diffusion model, consistently outperform their supervisors in both subject fidelity and prompt adherence. This is the first demonstration of such a phenomenon in the context of multimodal AR models for image generation.
@@ 5.1 Dataset and Evaluation
Dataset. Our experiments utilize the subject-driven image generation dataset introduced in DreamBooth [22]. We curate a diverse training dataset comprising 9 subjects: 4 live subjects (e.g., dog, cat) and 5 inanimate objects. Following DreamBooth [22], to ensure robust quantitative evaluation, we generate test images using 25 distinct prompts for each subject category. During testing, we produce 4 images per prompt for each subject, resulting in a comprehensive evaluation set of 225 images.
Evaluation Metrics. Subject-driven image generation requires evaluation along two critical dimensions: subject fidelity and prompt adherence. For subject fidelity, we employ two metrics: CLIP-I and DINO [3]. CLIP-I computes the cosine similarity between CLIP [20] image embeddings of real and generated images, while DINO calculates the same similarity using embeddings from ViT-S/16 [6]. For prompt fidelity, we use CLIP-T to measure the cosine similarity between CLIP image embeddings of generated images and text embeddings of their corresponding prompts.
@@ 5.2 Model Selection
Models. To validate the generalization capability of Proxy-Tuning across different model scales and architectures, we experiment with various diffusion supervisors and AR students for subject-driven image generation. For diffusion supervisors, we select models with different architectures and parameter scales but all trained using diffusion-based methods: SDXL [19] (U-Net, 2.6B parameters), SD3 Medium [7] (DiT, 2B parameters), SD3.5 Large [7] (DiT, 8B parameters), and FLUX.1 [dev] (DiT, 12B parameters). For AR students, we employ two models trained with autoregressive objectives: LlamaGen-XL from LlamaGen [24] (0.775B parameters) and FP-SFT@768 from Lumina-mGPT [15] (7B parameters). Training details are provided in Table 8 of the Appendix B. Unless otherwise specified, all experiments use parameter-efficient training with LoRA rather than full model fine-tuning.
@@ 5.3 Comparisons
Proxy-Tuning of AR Models. We evaluate our Proxy-Tuning method using various diffusion supervisors: SDXL (U-Net, 2.6B parameters), SD3 Medium (DiT, 2B parameters), SD3.5 Large (DiT, 8B parameters) and FLUX.1 [dev] (DiT, 12B parameters). Our approach first fine-tunes these diffusion models on specific targets and then uses their generated images to supervise the AR models. The results in Table 2, Table 3, and Figure 3 demonstrate that AR models not only effectively learn target characteristics but also surpass their diffusion supervisors in terms of CLIP-I and DINO metrics, regardless of the supervisor’s architecture. We term this surprising improvement the weak-to-strong generalization phenomenon.
Proxy-Tuning of Diffusion Models. To investigate whether the weak-to-strong generalization is unique to AR models, we apply Proxy-Tuning to diffusion models themselves. Specifically, this involves having diffusion supervisors supervise the fine-tuning process of diffusion models. By including more proxy subject samples and training iterations, our experiments in Table 4 consistently show that diffusion models perform worse than their supervisors in subject-driven image generation. This confirms that the weak-to-strong generalization phenomenon is specific to AR models.
@@
Why does the weak-to-strong generalization only occur in AR models but not in diffusion models? According to previous research [18, 35, 17], diffusion models often exhibit biased learning towards certain attributes in training datasets, and during inference, their progressive denoising process is susceptible to exposure bias[17], causing generated images to drift away from the target data distribution. These issues manifest themselves in subject-driven image generation tasks as diffusion models struggle to consistently reproduce subjects’ appearance attributes, leading to variance in generated images. In this case, when using Proxy-Tuning in diffusion models, these issues are exacerbated, leading to significant bias in the appearance of the subjects (as shown in Table 4). Compared to diffusion models, AR models encode the target subject’s image into a discrete token distribution, which is divided into two parts: a primary distribution composed of tokens representing the local appearance of the target subject, and a minor distribution formed by biased tokens introduced through diffusion models. During next-token prediction training, the AR model tends to fit the primary distribution (which captures the local appearance of the target subject) while filtering out the minor distribution formed by biased tokens. This allows the AR model to learn a more accurate representation of the target subject’s appearance.
@@ 5.4 Multiple Subjects Personalization
Given that AR models have demonstrated superior scalability to diffusion models [26], we investigate their capability in multi-subject personalization. While diffusion models typically require separate model instances or parameters for different subjects, we hypothesize that AR models can effectively learn multiple subjects in a single pass through our Proxy-Tuning framework. To test this hypothesis, we train a single AR model (LlamaGen) to simultaneously learn multiple subjects from the proxy data of separate diffusion supervisors. Each diffusion supervisor is first individually fine-tuned via LoRA to learn a specific subject. We then train the AR student to generate four randomly selected subjects using distinct prompts (denoted as $S^{*}_{1}$, $S^{*}_{2}$, etc.) designed to elicit subject-specific characteristics.
@@ 5.5 User Study
To complement our quantitative metrics and provide human perception-based evaluation, we conduct a user study comparing four approaches: direct subject tuning of diffusion and AR models, Proxy-Tuning of diffusion models, and our proposed Proxy-Tuning of AR models. Concretely, the diffusion model is SDXL and the AR model is Lumina-mGPT. Participants evaluate generated images based on three criteria: image quality, subject fidelity (how well the images preserve the subject’s characteristics), and prompt adherence (how well the images align with the given prompts). Further details of user study are shown in Section A.
@@ 5.6 Ablation Study
Impact of Proxy Images. To understand how the quantity of supervisor-generated images affects performance, we experiment with proxy datasets of varying sizes. As shown in left part of Figure 5, Proxy-Tuning demonstrates robust performance across different dataset sizes, suggesting that our method can maintain effectiveness even with a relatively small number of supervisor-generated images.
@@
Data Augmentation. A natural question is whether traditional data augmentation techniques could achieve similar benefits as our proxy-based approach. We compare Proxy-Tuning with geometric data augmentations (as shown in Table 7) applied to the original training images. While these augmentations preserve the target’s texture and color characteristics, Figure 6 shows that they lead to overfitting and limited semantic editability. This comparison reveals a key insight: AR models’ inherent capabilities in semantic understanding and generalization enable them to learn from the rich, contextual variations provided by diffusion supervisors, rather than just surface-level transformations. The superior performance of Proxy-Tuning over data augmentation suggests that AR models can effectively leverage their pre-trained knowledge of semantic relationships to extract and generalize subject-specific features, contributing to the weak-to-strong generalization phenomenon we observed. This aligns with AR models’ demonstrated advantages in scalability and semantic comprehension [29, 26].
Augmentations Strength
Random Horizontal Flipping 0.05
Random Vertical Flipping 0.05
Random Rotation (-5°, 5°)
Random Cropping (0.9, 1.0)
Table 7: Details of mild geometric data augmentations.
@@
Rank of LoRA. We investigate the sensitivity of Proxy-Tuning to different LoRA ranks. As shown in the right part of the Figure 5, our method maintains stable performance across various rank settings, consistently outperforming the subject tuning (LlamaGen) shown in Table 2 and 3. This robustness to LoRA rank selection further demonstrates the reliability of our approach in practical applications.
6 Conclusion

### v2

@@ 1 Introduction
Recently, multimodal autoregressive (AR) models [chern2024anole, team2024chameleon, liu2024lumina, ge2024seed, wang2024emu3], based on next-token prediction and transformer architecture [vaswani2017attention], have demonstrated remarkable capabilities across various multimodal tasks, including text-to-image (T2I) generation. These models have shown comparable performance to large diffusion models [rombach2022high, podell2023SDXL, esser2024scaling] in general image generation tasks. However, the application of AR models to subject-driven image generation, a task of significant user interest [ruiz2023dreambooth], remains relatively unexplored compared to the extensive research on diffusion models in this domain [ruiz2023dreambooth, chen2023disenbooth, alaluf2023neural, gal2022image, ye2023ip, li2024photomaker, wu2024infinite].
Subject-driven image generation [ruiz2023dreambooth, gal2022image, kumari2023multi] requires a model to learn the visual appearance of a specific subject from a small set of reference images and then generate novel images of that subject based on user prompts. While diffusion models [ho2020denoising, rombach2022high] have been successfully adapted for this task using methods like DreamBooth [ruiz2023dreambooth], our initial attempts to tailor AR models for subject-driven generation revealed significant challenges.
We explore two primary approaches to adapt AR models for this task: parameter-efficient fine-tuning using LoRA [hu2021lora] and end-to-end fine-tuning. In both cases, we bound the given subject to a pre-defined token (e.g., S*) for use during inference. However, our experiments uncover limitations in both approaches. LoRA-based fine-tuning struggles to capture the specific subject’s appearance accurately, while end-to-end fine-tuning leads to diminished text fidelity in the generated images and compromises the model’s original capabilities.
To address these limitations, we introduce Proxy-Tuning, a novel method that leverages diffusion models to enhance AR models’ capabilities in subject-specific image generation. Our approach involves first fine-tuning a relatively weaker diffusion model on the reference images of the specific subject, and then uses the text-image pairs generated by this fine-tuned diffusion model to supervise the fine-tuning of more powerful AR models.
Surprisingly, we discover a striking weak-to-strong generalization [burns2023weak] phenomenon: fine-tuned AR models consistently outperform their diffusion model supervisors in both subject fidelity and prompt adherence. This trend holds true across different combinations and architectures of teacher (diffusion) and student (AR) models. Further analysis reveals scenarios where AR models particularly excel, e.g., multi-subject compositions and contextual understanding.
@@ 2.1 Subject-Driven Image Generation
Subject-driven image generation is a task that focuses on generating images related to a specific subject, where the model learns from a few reference images of the subject and combines with given prompts to create novel images of that subject. Currently, the existing methods regarding subject-driven image generation are mainly divided into two categories. The first category is tuning-based methods [ruiz2023dreambooth, chen2023disenbooth, alaluf2023neural, gal2022image, kumari2023multi], which require fine-tuning the model for each specific target. Specifically, DreamBooth [ruiz2023dreambooth] binds the given subject to a specially pre-defined token (e.g., S*) and introduces a class-specific prior preservation loss to preserve the original knowledge of the pretrained diffusion model. Disenbooth [chen2023disenbooth] proposes to enhance the disentanglement of the specific subject and background features during the generation process, aiming to improve the quality and flexibility of the generated images. NeTI [alaluf2023neural] focuses on improving the neural network architectures used in subject-driven image generation to achieve better visual fidelity and semantic consistency. The another category is tuning-free methods [ye2023ip, li2024photomaker, wu2024infinite, xiao2024fastcomposer, wang2024instantid, ma2024subject, shi2024instantbooth, wei2023elite]. Among these methods, IP-Adapter [ye2023ip] introduces a decoupled cross-attention mechanism to capture and fuse the image embedding input and text embedding input to guide the process of image generation. PhotoMaker [li2024photomaker] extracts image features by using a pretrained CLIP image encoder and projects the image features to the text latent space through a lightweight mapper. Eventually, these image features are concatenated with the text features and then input into the model. Similarly, ELITE [wei2023elite] also extracts image features by using a pretrained encoder, and incorporates the image features into the text latent space and the cross-attention module respectively by means of global mapping and local mapping.
@@ 2.2 Multimodal Autoregressive Models
Multimodal autoregressive (AR) models [sun2024autoregressive, team2024chameleon, liu2024lumina, wang2024emu3, ge2024seed, chern2024anole], based on next-token prediction and transformer architecture, have shown impressive abilities in various multimodal tasks, especially in the realm of text-to-image (T2I) generation. LlamaGen [sun2024autoregressive] explores the application of the ”next-token prediction” paradigm from large language models to visual generation. By scaling the widely used autoregressive Llama backbones, it attains excellent image generation performance. Similarly, Chameleon [team2024chameleon] also adopts the autoregressive generation approach, capable of transforming images and text into discrete tokens for processing, showing excellent performances in various tasks such as visual question answering and image generation. Based on Chameleon, Lumina-mGPT [liu2024lumina] enhances the training datasets and achieves flexible text-to-image generation and controllable generation, also capable of handling vision recognition tasks. Moreover, Emu3 [wang2024emu3] further includes the video datasets in the training datasets and trains the transformer model only with next-token prediction by encoding images, texts and videos into a discrete space and jointly training on multimodal mixed sequences.
3 Analysis: Subject-Driven AR Tuning
@@
LoRA SDXL [podell2023SDXL] 0.8002 0.3225 0.7272
End-to-End Lumina-mGPT [liu2024lumina] 0.6974 0.2652 0.5338
LoRA Lumina-mGPT [liu2024lumina] 0.6752 0.2956 0.5088
Table 1: Quantitative evaluation of DreamBooth [ruiz2023dreambooth] subject-driven fine-tuning on AR model. Details about the metrics are provided in Section 5.1.
Figure 2: Framework of Proxy-Tuning. Provided with a limited number of images of a particular target, we initially conduct subject-tuning (i.e., DreamBooth) on the diffusion model. Then the diffusion model is employed as a supervisor to supervise the fine-tuning of the AR model. Direct subject-tuning of the AR model results in an unsatisfactory acquisition of the target appearance, and the utilization of data augmentations in subject-tuning manifests a deficiency in semantic editability (as discussed in Section 5.6). Conversely, Proxy-Tuning effectively captures the target appearance and simultaneously showcases excellent semantic editability.
To explore subject-driven image generation capabilities in AR models, we initially explored the settings in DreamBooth [ruiz2023dreambooth] of fine-tuning models with few subject images. Our investigation focused on two primary scenarios: end-to-end fine-tuning of the entire model and parameter-efficient tuning using LoRA [hu2021lora]. Following the DreamBooth paradigm, we associate the target subject with a designated token (e.g., ”S*”) for inference purposes.
@@
Notably, these challenges are less prominent in diffusion models under similar few-shot subject-driven settings [ruiz2023dreambooth]. This disparity can be attributed to the inherent characteristics of AR models: they generate images through sequential token predictions where each token heavily depends on previous ones. This autoregressive nature makes them particularly sensitive to parameter updates when trained on very few subject images, as even slight perturbation in token prediction can propagate and amplify throughout the generation sequence. In contrast, diffusion models’ parallel denoising process demonstrates greater robustness to few-shot fine-tuning, better preserving both subject fidelity and semantic consistency.
@@ 4 Method: Proxy-Tuning
Notably, we discover a compelling weak-to-strong phenomenon in image generation, similar to observations in natural language processing [burns2023weak]. Our findings show that AR models, when trained under the supervision of a weak diffusion model, consistently outperform their supervisors in both subject fidelity and prompt adherence. This is the first demonstration of such a phenomenon in the context of multimodal AR models for image generation.
@@ 5.1 Dataset and Evaluation
Dataset. Our experiments utilize the subject-driven image generation dataset introduced in DreamBooth [ruiz2023dreambooth]. We curate a diverse training dataset comprising 9 subjects: 4 live subjects (e.g., dog, cat) and 5 inanimate objects. Following DreamBooth [ruiz2023dreambooth], to ensure robust quantitative evaluation, we generate test images using 25 distinct prompts for each subject category. During testing, we produce 4 images per prompt for each subject, resulting in a comprehensive evaluation set of 225 images.
Evaluation Metrics. Subject-driven image generation requires evaluation along two critical dimensions: subject fidelity and prompt adherence. For subject fidelity, we employ two metrics: CLIP-I and DINO [caron2021emerging]. CLIP-I computes the cosine similarity between CLIP [radford2021learning] image embeddings of real and generated images, while DINO calculates the same similarity using embeddings from ViT-S/16 [dosovitskiy2020image]. For prompt fidelity, we use CLIP-T to measure the cosine similarity between CLIP image embeddings of generated images and text embeddings of their corresponding prompts.
@@ 5.2 Model Selection
Models. To validate the generalization capability of Proxy-Tuning across different model scales and architectures, we experiment with various diffusion supervisors and AR students for subject-driven image generation. For diffusion supervisors, we select models with different architectures and parameter scales but all trained using diffusion-based methods: SDXL [podell2023SDXL] (U-Net, 2.6B parameters), SD3 Medium [esser2024scaling] (DiT, 2B parameters), SD3.5 Large [esser2024scaling] (DiT, 8B parameters), and FLUX.1 [dev] (DiT, 12B parameters). For AR students, we employ two models trained with autoregressive objectives: LlamaGen-XL from LlamaGen [sun2024autoregressive] (0.775B parameters) and FP-SFT@768 from Lumina-mGPT [liu2024lumina] (7B parameters). Unless otherwise specified, all experiments use parameter-efficient training with LoRA rather than full model fine-tuning.
@@ 5.3 Comparisons
Proxy-Tuning of AR Models. We evaluate our Proxy-Tuning method using various diffusion supervisors: SDXL (U-Net, 2.6B parameters), SD3 Medium (DiT, 2B parameters), SD3.5 Large (DiT, 8B parameters), and FLUX.1 [dev] (DiT, 12B parameters). Our approach first fine-tunes these diffusion models on specific targets, then uses their generated images to supervise the AR models. The results in Table 2, Table 3, and Figure 3 demonstrate that AR models not only effectively learn target characteristics but also surpass their diffusion supervisors in terms of CLIP-I and DINO metrics, regardless of the supervisor’s architecture. We term this surprising improvement the weak-to-strong generalization phenomenon.
Proxy-Tuning of Diffusion Models. To investigate whether the weak-to-strong generalization is unique to AR models, we apply Proxy-Tuning to diffusion models. Specifically, this involves having diffusion supervisors supervise the fine-tuning process of diffusion models. By including more proxy subject samples and training iterations, experiments in Table 4 consistently show that diffusion models perform worse than their supervisors in subject-driven image generation, which confirms that the weak-to-strong generalization phenomenon is specific to AR models.
@@
Why does the weak-to-strong generalization only occur in AR models but not in diffusion models? According to previous research [parihar2024balancing, yue2024few, ning2023elucidating], diffusion models often exhibit biased learning towards certain attributes in training datasets, and during inference, their progressive denoising process is susceptible to exposure bias[ning2023elucidating], causing generated images to drift away from the target data distribution. These issues manifest themselves in subject-driven image generation tasks as diffusion models struggle to consistently reproduce subjects’ appearance attributes, leading to variance in generated images. In this case, when using Proxy-Tuning in diffusion models, these issues are exacerbated, leading to significant bias in the appearance of the subjects (as shown in Table 4). Compared to diffusion models, AR models encode the target subject’s image into a discrete token distribution, which is divided into two parts: a primary distribution composed of tokens representing the local appearance of the target subject, and a minor distribution formed by biased tokens introduced through diffusion models. During next-token prediction training, the AR model tends to fit the primary distribution (which captures the local appearance of the target subject) while filtering out the minor distribution formed by biased tokens. This allows the AR model to learn a more accurate representation of the target subject’s appearance.
@@ 5.4 Multiple Subjects Personalization
Given that AR models have demonstrated superior scalability to diffusion models [tian2024visual], we investigate their capability in multi-subject personalization. While diffusion models typically require separate model instances or parameters for different subjects, we hypothesize that AR models can effectively learn multiple subjects in a single pass through our Proxy-Tuning framework.
To test this hypothesis, we train a single AR model (LlamaGen) to simultaneously learn multiple subjects from the proxy data of separate diffusion supervisors. Each diffusion supervisor is first individually fine-tuned via LoRA to learn a specific subject. We then train the AR student to generate four randomly selected subjects using distinct prompts (denoted as $S^{*}_{1}$, $S^{*}_{2}$, etc.) designed to elicit subject-specific characteristics.
@@ 5.5 User Study
To complement our quantitative metrics and provide human perception-based evaluation, we conduct a user study comparing four approaches: direct subject tuning of diffusion and AR models, Proxy-Tuning of diffusion models, and our proposed Proxy-Tuning of AR models. Concretely, the diffusion model is SDXL and the AR model is Lumina-mGPT. Participants evaluate generated images based on three criteria: image quality, subject fidelity (how well the images preserve the subject’s characteristics), and prompt adherence (how well the images align with the given prompts). The study involves 9 participants. The participants are asked to grade the subject images based on three criteria: image quality, subject fidelity, and prompt fidelity. A comprehensive grading is required for the several groups of images of the four methods in the context of three cases. Each case is composed of a subject image and several groups of images obtained from four methods. In each group, there is one piece of text and two images. The results are rated on a scale of 1 to 5, with 5 being the highest. We include direct subject tuning of diffusion and AR models, Proxy-Tuning of diffusion models, and our proposed Proxy-Tuning of AR models.
@@ 5.6 Ablation Study
Impact of Proxy Images. To understand how the quantity of supervisor-generated images affects performance, we experiment with proxy datasets of varying sizes. As shown in Figure 6, Proxy-Tuning demonstrates robust performance across different dataset sizes, suggesting that our method can maintain effectiveness even with a relatively small number of supervisor-generated images.
@@
Data Augmentation. A natural question is whether data augmentation techniques could achieve similar benefits as our method. We compare Proxy-Tuning with geometric data augmentations applied to the original training images. Specifically, the geometric data augmentations include random horizontal and vertical flipping with a strength of 0.05, random rotation with a strength from -5°to 5°, and random cropping with a strength from 0.9 to 1.0. While these augmentations preserve the target’s texture and color characteristics, Figure 5 shows that they lead to overfitting and limited semantic editability. This comparison reveals a key insight: AR models’ inherent capabilities in semantic understanding and generalization enable them to learn from the rich, contextual variations provided by diffusion supervisors, rather than just surface-level transformations. The superior performance of Proxy-Tuning over data augmentation suggests that AR models can effectively leverage their pre-trained knowledge of semantic relationships to extract and generalize subject-specific features, contributing to the weak-to-strong generalization phenomenon we observed.
@@
Rank of LoRA. We investigate the sensitivity of Proxy-Tuning to different LoRA ranks. As shown in Figure 6, our method maintains stable performance across various rank settings, consistently outperforming the subject tuning (LlamaGen) shown in Table 2 and 3. This robustness to LoRA rank selection further demonstrates the reliability of our approach in practical applications.
6 Conclusion and Limitations